OpenStudy (anonymous):

√(x) - 1 ÷ x^2-1 limit ->1

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{\sqrt{x}-1}{x^2-1} ?\]

OpenStudy (anonymous):

yeah, thats it!

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{\sqrt{x}-1}{(x-1)(x+1)}=\lim_{x \rightarrow 1}\frac{\sqrt{x}-1}{(\sqrt{x}-1)(\sqrt{x}+1)(x+1)}\]

OpenStudy (anonymous):

I applied Hopital's Rule and got 1/4 which is the same as myininaya.

myininaya (myininaya):

\[=\lim_{x \rightarrow 1}\frac{1}{(\sqrt{x}+1)(x+1)}=\frac{1}{(\sqrt{1}+1)(1+1)}=\frac{1}{2(2)}=\frac{1}{4}\]

myininaya (myininaya):

yes you can do l'hopital too! :)

OpenStudy (anonymous):

\[(1/2x ^{-1/2})/(2x)\] \[1/(4x ^{3/2})\]

myininaya (myininaya):

i factored x-1 as (sqrt{x}-1)(sqrt{x}+1)

OpenStudy (anonymous):

Hopital's only works since the top and bottom are 0 when eveluated at 1.

OpenStudy (anonymous):

Nice factoring btw :)

myininaya (myininaya):

thanks nice "l'hopitaling" there

myininaya (myininaya):

if you didn't like me factoring that we could had done this alittle differently \[\lim_{x \rightarrow 1}\frac{\sqrt{x}-1}{x^2-1}*\frac{\sqrt{x}+1}{\sqrt{x}+1}=\lim_{x \rightarrow 1}\frac{x-1}{(\sqrt{x}+1)(x^2-1)}\] \[=\lim_{x \rightarrow 1}\frac{(x-1)}{(\sqrt{x}+1)(x-1)(x+1)}=\lim_{x \rightarrow 1}\frac{1}{(\sqrt{x}+1)(x+1)}\]